5 It 5 df Pe s a te bu 7; 9, c3 of th ahPaG 7 r ti uddG crr aae p) Pte eatrePlease only focus on requirement # 2 & show ALL work
Answer #1

Scenario (a):-

Total Units = 6500 pounds

Sales Value = 6500*2.4 = 15600

Total Cost = 6500*2.3 = 14950

Saving in processing deparment due to discontinuation of Silpol = 1500

Operating income = 15600-14950+1500 = 2150

Scenario (b):-

Units of Silpol = 3500

For 1unit of Silpol 1/4 pound of GR37 required, So for 3500 units of Silpol, 875 pounds (3500*1/4) of GR37 requireed

Remaining Pounds after 3500 units of Silpol = 6500-875 = 5625

Particulars GR37 Silpol
Sales 5625*2.4 = 13500 3500*5 = 17500
Less: Costs 5625*2.3 = 12937.50
Additional processing cost 3500*4 = 14000
Variable Selling Cost 3500*0.20 = 700
Operating Income 562.50 2800

Total Operating Income = 562.50+2800 = 3362.50

Scenario (c):-

Units of Silpol = 7500

Required pounds of GR37 = 7500*1/4=1875

Balance pounds for GR37 = 6500-1875 = 4625

Particulars GR37 Silpol
Sales 4625*2.4 = 11100 7500*5 = 37500
Less: Costs 4625*2.3 = 10637.50
Additional processing cost 7500*4 = 30000
Variable Selling Cost 7500*0.20 = 1500
Operating Income 462.50 6000

Total Operating Income = 462.50+6000 = 6462.50

Scenario (d):-

Units of Silpol = 9500

Required pounds of GR37 = 9500*1/4 = 2375

Balance pounds for GR37 = 6500-2375 = 4125

Particulars GR37 Silpol
Sales 4125*2.4 = 9900 9500*5 = 47500
Less: Costs 4125*2.3 = 9487.50
Additional processing cost 9500*4 = 38000
Variable Selling Cost 9500*0.20 = 1900
Operating Income 412.50 7600

Total Operating Income = 412.50+7600 = 8012.50